# Generic motives and motivic cohomology of fields

### v1

@@ 4.2. Positive genus curves
splits and gives a canonical isomorphism $\operatorname{M}^{c}(C)=\operatorname{coKer}(\mathbbm{1}\rightarrow\operatorname{M}(C))=\underline{A}\oplus\mathbbm{1}(1)[2]$. Therefore, according to (4.2.1.a), one gets an isomorphism
@@
One deduces the desired isomorphism as the factor $\mathbbm{1}$ is cut-off by the projection $p:C\rightarrow\operatorname{\mathrm{Spec}}(k)$. ∎
@@
Recall that $\mathbbm{1}(1)[1]=\mathbb{G}_{m}$. Therefore, the preceding formula reflects the duality of abelian variety. Note in particular that, as $A$ is the Albanese variety of a pointed smooth projective curve, it admits a polarization which explains the preceding formula. For a general abelian variety $A$, one would expect $\underline{A}^{\vee}=\underline{(A^{*})}(-1)[-2]$ where $A^{*}$ is the dual abelian variety. This formula is proved for étale motives (and therefore for $\operatorname{DM}(k,\mathbb{Q})$) by Kahn and Barbieri-Viale in [BVK16, 4.5.3].
@@ 4.3. Higher dimension
Given the preliminary sections, the reader will not be surprised by the fact generic motives are gigantic in general. We can at least illustrate the first assertion by the following fact which partially extends 4.1.5.
@@
Let $\varphi:K\rightarrow L$ be an extension of function fields over $k$ of transcendence degree $n>0$. To facilitate the notation, we consider the pro-motive $T=\underset{x\in\mathbb{A}^{1}_{K,(0)}}{\text{\rm``$\prod$''}}\operatorname{\hat{M}}(\kappa_{x})\{1\}$.
Consider a transcendental basis $(t_{1},...,t_{n})$ of $L/K$, and let $K_{i}=K(t_{1},...,t_{i})\subset L$. We will consider the finite morphism $\psi:K_{n}\rightarrow L$ and for any $i\leq j$, we let $\varphi_{i,j}:K_{i}\rightarrow K_{j}$ be the obvious inclusion.
@@
where one defines the projector $\pi_{i}^{n}:\operatorname{\hat{M}}(K_{n})\rightarrow T\{i\}$ inductively using the formulas:
@@
where $\phi:\operatorname{\hat{M}}(K_{1})\rightarrow T\{1\}$ is the projection map from 4.1.2, and $\partial_{n}:\operatorname{\hat{M}}(K_{n})\rightarrow\operatorname{\hat{M}}(K_{n-1})\{1\}$ is the residue map coming from the valuation on $K_{n}$ associated with the variable $t_{n}$.
@@
Using 4.1.4 (or the degree formula directly), $\psi^{*}$ is a split epimorphism, so that on reduces to the purely transcendental case $L=K_{n}$. Then it follows by induction on $n>0$: the case $n=1$ is precisely 4.1.2. Then we get a composite epimorphism:
$\operatorname{\hat{M}}(K_{n})$ $\rightarrow\operatorname{\hat{M}}(K_{n-1})\oplus\operatorname{\hat{M}}(K_{n-1})\{1\}$ 4.1.2 to $K_{/}K_{n-1}$
@@
and we conclude by induction using the $(n-1)$-case (up to discarding the first factor). The formulas for the projectors follow from this induction. ∎
@@
The preceding proposition only gives a rough “lower bound” on the different pro-motives that a given generic contains. Giving the analysis on the preceding section, formula 4.1.2 does not allow to compute the motive of even a purely transcendental extension of degree more than $1$. On the other hand, it shows that generic motives of transcental extensions of degree $n$ are usually non-trivial extensions of all motives of dimension less than $n$.
4.3.3.
Surfaces. We will illustrate the preceding remark in transcendental degree $2$. Let $S$ be a geometrically connected smooth projective surface over a field $k$ of characteristic 00. We let $K$ be its function field.
@@
which we will view in the homotopy category $\operatorname{Ho}\operatorname{DM}(k,\mathbb{Q})$. Moreover, according to [KMP07], this decomposition can be refined: the component $\operatorname{h}^{2}(S)$ can be split into:
@@
respectively the algebraic and transcendental part of $\operatorname{h}^{2}$. Note that by definition, the algebraic part is of the form: $\operatorname{h}^{2}_{alg}=b_{1}.\mathbbm{1}(-1)[-2]$ where $b_{1}$ is the first Betti number of the surface $S$. Beware that in loc. cit., the motive $\operatorname{h}^{2}_{tr}(S)$ is denoted by $t_{2}(S)$. Recall that it is shown in Corollary 14.8.11 of loc. cit. that the functor $S\mapsto\operatorname{h}^{2}_{tr}(S)$ is a birational invariant of $S$ [^26] As recalled to me by Jan Nagel, this property can also be shown directly by using the blow-up formula for Chow motives combined with the fact any birational morphism of surfaces in characteristic 00 can be written as a sequence of blow-ups of closed points. In particular, it depends only on the function field $K$ of $S$.
Dually, we get a decomposition of the homological motive, in $\operatorname{DM}(k,\mathbb{Q})$:
@@
Letting $A$ be the Albanese scheme associated with $S$, we get an identification $\operatorname{M}_{1}(S)=\underline{A}$, and by duality: $\operatorname{M}_{3}(S)=\underline{A}^{\vee}(2)[4]\simeq\underline{A}(1)[2]$ (apply 4.2.6). Let us sumarize the above discussion.
@@
$\operatorname{M}(S)=\mathbbm{1}\oplus\underline{A}\oplus b_{1}.\mathbbm{1}(1)[2]\oplus\operatorname{M}_{2}^{tr}(S)\oplus\underline{A}(1)[2]\oplus\mathbbm{1}(2)[4].$
Remark 4.3.5.
We can view the above decomposition in the effective category $\operatorname{DM^{eff}}(k,\mathbb{Q})$ of motivic complexes. Then it is remarkable that, apart the mysterious part $\operatorname{M}_{2}^{tr}(S)$ for which we do not know, all the components of the decomposition lives in homotopical degree $[0,\dim(S)]$.
4.3.6.
To go further, we will need to use the coniveau filtration viewed internally inside pro-motives, as in [Dég12] or [DP24]. [^27] We follow the notation of the latter which seems better suited. We still consider a geometrically smooth projective surface $S$ over $k$, with function field $K$.
Recall that a flag in $S$ is a decreasing sequence $(Z^{p})_{p\geq 0}$ of closed subschemes of $S$ such that $\operatorname{codim}_{S}(Z^{p})\geq p$. The set $\mathscr{F}lag(S)$ of flags of $S$, ordered by term-wise inclusion, is filtered. We define an increasing filtration of $S$ by pro-schemes, the coniveau filtration, by setting:
$S^{\leq p}:=\underset{Z^{*}\in\mathscr{F}lag(S)}{\text{\rm``lim''}}S-Z^{p+1}.$
By applying the functor $\operatorname{M}$ term-wise, we then get a canonical homotopy exact sequence in $\operatorname{pro-\mskip-3.0mu}\operatorname{DM_{gm}}(k)$:
@@
Then one can extend the computation of 4.2.2 in the transcendental $2$ case as follows.
@@
Consider the above notation. Then there exists homotopy exact sequences in $\operatorname{pro-\mskip-3.0mu}\operatorname{DM_{gm}}(k,\mathbb{Q})$:
@@
$\underset{s\in S^{(2)}}{\text{\rm``$\prod$''}}\operatorname{\hat{M}}(\kappa_{s})\{2\}\rightarrow\operatorname{\hat{M}}(S^{\leq 1})\rightarrow\mathbbm{1}\oplus\underline{A}\oplus b_{1}.\mathbbm{1}(1)[2]\oplus\operatorname{M}_{2}^{tr}(S)\oplus\underline{A}(1)[2]\oplus\mathbbm{1}(2)[4]$
where the $\partial_{x}$ is the residue map (see 3.3.5(D4*)) associated with the valuation on $K$ corresponding to the divisor with generic point $x\in S^{(1)}$.
In particular, the motive of the transcendental degree $2$ function field $K/k$ can be computed in terms of a bi-extension of generic motives associated with fields of lower transcendental degrees on the one hand, and of pure motives, including the transcendental motive of $S$, on the other hand.
@@
This follows immediately from the exact sequences (4.3.6.a) and the preceding discussion, once one notices that $\operatorname{M}(S^{\leq 2})=\operatorname{M}(S)$ and $\operatorname{M}(S^{\leq 0})=\operatorname{M}(S^{=0})=\operatorname{\hat{M}}(K)$. ∎
@@ 5. Motivic cohomology of fields
The structure of the generic motive of a function field $F/k$ is reflected in motivic cohomology, as follows from the formula:
(5.0.1.a) $\operatorname{H}_{\mathrm{M}}^{n,i}(K)=\operatorname{Hom}_{\operatorname{pro-\mskip-3.0mu}\operatorname{DM}(k)}(\operatorname{\hat{M}}(K),\mathbbm{1}(i)[n]).$
@@
The latter group as several explicit presentation. For the non trivial case $i>0$, we consider the non positively graded homological complex made of free abelian groups
@@
where all schemes and products are taken implicitly over the field $K$, $\mathbb{G}_{m}^{\wedge,i}$ is the $i$-th smash product of the pointed scheme $(\mathbb{G}_{m},1)$ (see also 2.2.1 for this notation). Thus the group $c_{K}(\Delta^{n},\mathbb{G}_{m}^{\wedge,i})$ is the free abelian group generated by integral closed subschemes $Z\subset\Delta^{n}\times\mathbb{G}_{m}^{\wedge,i}$ which are finite dominant over $\Delta^{n}_{K}$ and which does not belong to a closed subscheme of the form $\Delta^{n}\times(\mathbb{G}_{m}^{\wedge,j}\times\{1\}\times\mathbb{G}_{m}^{\wedge,i-j-1})$. The differentials are given by the alternate sum of the operation of taking intersection along faces $\Delta^{n-1}_{i}\times\mathbb{G}_{m}^{\wedge,n}$, operations which are always well-defined and can be computed using Serre’s Tor formula.
Then one gets the following formula for motivic cohomology of $K$
@@
As both abelian groups $\operatorname{Ker}(d_{i-n})$ and $\operatorname{Im}(d_{i-n-1})$ are free, this is therefore an explicit presentation of motivic cohomology. Of course, it is useless for practical computations, but it gives at least a bound for the size of motivic cohomology of fields (see 5.2.4 for more). [^28] One gets a similar presentation by considering Bloch’s higher Chow groups.
@@ 5.1. Borel’s computations and $\lambda$-structures
In this subsection, we recall the following fundamental computation of motivic cohomology, mainly due to Borel if one neglects weights, which were determined by Beilinson using his not less fundamental theory of (higher) regulators. The main argument of Beilinson was explained (and corrected!) with all details by Burgos in [BG02].
@@
Let $K$ be a number field. Let $r_{1}$ and $r_{2}$ be respectively the number of real and complex embeddings of $K$. Then for any couple of integers $(n,i)$, one has:
@@
$\operatorname{H}_{\mathcyr{B}}^{n,i}(\mathcal{O}_{K})=\begin{cases}\mathbb{Q}&n=i=0,\\ \mathcal{O}_{K}^{*}\otimes_{\mathbb{Z}}\mathbb{Q}&n=i=1,\\ \mathbb{Q}^{r_{2}}&n=1,i>1,i\text{ even},\\ \mathbb{Q}^{r_{1}+r_{2}}&n=1,i>1,i\text{ odd},\\ \operatorname{Pic}(O_{K})_{\mathbb{Q}}&n=2,i=1,\\ 0&\text{otherwise.}\end{cases}$
We will give a proof of this theorem, or rather a reduction to Borel’s computation of the stable real cohomology of arithmetic groups ([Bor74]). We first recall the arguments in low degree for the comfort of the reader (the latter are now well established facts).
The computations for $i\leq 1$ follow from (2.2.1.b). The same holds for the vanishing $\operatorname{H}_{\mathcyr{B}}^{n,i}(K)$ (resp. $\operatorname{H}_{\mathcyr{B}}^{n,i}(\mathcal{O}_{K})$) for $n-i>0$ (resp. $n-i>1$). We can consider the coniveau spectral sequences for the $1$-dimensional regular scheme $X=\operatorname{Spec}(\mathcal{O}_{K})$ with coefficients in $n$-th twisted rational motivic cohomology and rational K-theory: [^29] The second one is a direct sum of instances of the first one, but it is practical to use both forms.
@@
[…]
@@
Note in particular that $\operatorname{H}_{\mathcyr{B}}^{3,2}(\mathcal{O}_{K})$ is necessarily 00 according to the first isomorphism. As a summary, we have obtained the lower $K$-groups as follows:
@@
Thus, we are reduced to compute $K_{n}(\mathcal{O}_{K})_{\mathbb{Q}}$ for $n\geq 2$, and to determine its graduation for the $\gamma$-filtration to get the above calculation. As already mentioned, the first task was achieved by Borel, in his famous paper [Bor74]. The second task follows from the comparison of the Borel and the Beilinson regulators, as made explicit in [BG02] (see also [Rap88] for an earlier account). In the remaining of this section, we will explicitly recall the computations of Borel and show how to give a direct argument for the determination of the $\gamma$-filtration.
@@
Rational K-theory and homology of groups. [^30] These facts are well-known, but we still recall them for completeness, and for introducing our notation. Given any commutative ring $A$ and an integer $n>1$, one gets the following classical computation:
@@
where $\operatorname{GL}(A)$ (resp. $\operatorname{SL}(A)$) is the infinite general (resp. special) linear group associated with $A$, and $X^{+}$ is Quillen $+$-construction applied to a space $X$ (see also 7.0.2). The first isomorphism is Quillen’s comparison of the plus and Q constructions ([Gra76], [God] with more details). To get the second isomorphism one considers the obvious exact sequence of groups:
@@
After applying the classifying space construction, followed by the plus construction, one gets a fibration sequence in the homotopy category: [^31] Note that the $+$-construction respects fibration sequences, or equivalently homotopy fibers, as it is a localization functor.
$\operatorname{BSL}(A)^{+}\rightarrow\operatorname{BGL}(A)^{+}\rightarrow\mathrm{B}(A^{\times}).$
Thus the associated exact sequence of homotopy groups allows to conclude as $A^{\times}$ is a finite abelian group.
Next we recall that $\operatorname{BSL}(A)^{+}$ is an h-space323232in fact, the +=Q theorem of Quillen shows it is a loop space., therefore the Milnor-Moore/Cartan-Serre theorem implies that the Hurewicz map
$\pi_{n}(\operatorname{BSL}(A)^{+})\rightarrow H_{n}^{sing}(\operatorname{BSL}(A),\mathbb{Z})$
@@
$H_{*}(SL(A),\mathbb{Q})=H_{*}^{sing}(\operatorname{BSL}(A),\mathbb{Q})\simeq H_{*}^{sing}(\operatorname{BSL}(A)^{+},\mathbb{Q}).$
@@
We next fix an integer $r>0$, consider the arithmetic group $\Gamma_{r}=\operatorname{SL}_{r}(\mathcal{O}_{K})$, which can be seen as a subgroup of the real Lie group [^33] The latter can also be seen as the real points of the algebraic $\mathbb{Q}$-group scheme obtained by Weil scalar restriction of the $K$-group scheme $\operatorname{SL}_{r,K}$.
@@
We let $\nu_{r}:\Gamma_{r}\rightarrow G_{r}$ be the natural inclusion, where $\Gamma_{r}$ is seen with the discrete topology. Viewing $\mathbb{R}$ as a discrete module over $G_{r}$ (resp. $\Gamma_{r}$), one can define the continuous cohomology $H^{*}_{ct}(-,\mathbb{R})$ with coefficient in $\mathbb{R}$ (see e.g. [BW00, IX, §1]). As an application of the main result of [Bor74], one gets the following fundamental computation.

### v3

@@ 4.3. Curves of positive genus
splits, giving a canonical isomorphism $\operatorname{M}^{c}(C)=\operatorname{coKer}(\mathbbm{1}\rightarrow\operatorname{M}(C))=\underline{A}\oplus\mathbbm{1}(1)[2]$. Thus, according to (4.3.1.a), one gets an isomorphism
@@
One deduces the desired isomorphism as the factor $\mathbbm{1}$ is eliminated by the projection $p:C\rightarrow\operatorname{\mathrm{Spec}}(k)$. ∎
@@
Recall that $\mathbbm{1}(1)[1]=\mathbb{G}_{m}$. Therefore, the preceding formula reflects the duality of abelian varieties. Note in particular that, as $A$ is the Albanese variety of a pointed smooth projective curve, it admits a polarization which explains the preceding formula. For a general abelian variety $A$, one would expect $\underline{A}^{\vee}=\underline{(A^{*})}(-1)[-2]$ where $A^{*}$ is the dual abelian variety. This formula was proved for étale motives (and consequently for $\operatorname{DM}(k,\mathbb{Q})$) by Kahn and Barbieri-Viale in [BVK16, 4.5.3].
@@ 4.4. Higher dimension
Given the preliminary sections, the reader will not be surprised that generic motives are gigantic in general. We can at least illustrate this assertion by the following result which partially extends 4.2.5.
@@
Let $\varphi:K\rightarrow L$ be an extension of function fields over $k$ of transcendence degree $n>0$. For simplicity, we consider the pro-motive $T=\underset{x\in\mathbb{A}^{1}_{K,(0)}}{\text{\rm``$\prod$''}}\operatorname{\hat{M}}(\kappa_{x})\{1\}$.
Fix a transcendental basis $(t_{1},...,t_{n})$ of $L/K$, and let $K_{i}=K(t_{1},...,t_{i})\subset L$. We consider the finite morphism $\psi:K_{n}\rightarrow L$ and for any $i\leq j$, we let $\varphi_{i,j}:K_{i}\rightarrow K_{j}$ be the obvious inclusion.
@@
where the projector $\pi_{i}^{n}:\operatorname{\hat{M}}(K_{n})\rightarrow T\{i\}$ is defined inductively using the formulas:
@@
where $\phi:\operatorname{\hat{M}}(K_{1})\rightarrow T\{1\}$ is the projection map from 4.2.2, and $\partial_{n}:\operatorname{\hat{M}}(K_{n})\rightarrow\operatorname{\hat{M}}(K_{n-1})\{1\}$ is the residue map arising from the valuation on $K_{n}$ associated with the variable $t_{n}$.
@@
Using 4.2.4 (or the degree formula directly), $\psi^{*}$ is a split epimorphism, so that we are reduced to the purely transcendental case $L=K_{n}$. Then it follows by induction on $n>0$: the case $n=1$ is precisely 4.2.2. Then we get a composite epimorphism:
$\operatorname{\hat{M}}(K_{n})$ $\rightarrow\operatorname{\hat{M}}(K_{n-1})\oplus\operatorname{\hat{M}}(K_{n-1})\{1\}$ 4.2.2 for $K_{n}/K_{n-1}$
@@
and we conclude by induction using the $(n-1)$-case (up to discarding the first factor). The formulas for the projectors follow from the induction. ∎
@@
The preceding proposition provides only a preliminary decomposition of the various pro-motives contained within a given generic motive. Despite the analysis in the preceding section, the formula in 4.2.2 is not sufficient to explicitly compute the structure of the motive of a purely transcendental extension of degree greater than one. Nevertheless, it reveals that generic motives of transcendence degree $n$ typically form non-trivial extensions of lower-dimensional motives. In particular, these motives are not mere direct sums or simple products of lower-dimensional components but involve intricate extension structures, reflecting the increasing complexity of their cohomological and motivic information.
4.4.3.
Surfaces. We now assume that $k$ is algebraically closed. We will illustrate the preceding remark in transcendence degree $2$. Let $S$ be a geometrically connected smooth projective surface over $k$. We let $K$ be its function field.
@@
which we will consider in the homotopy category $\operatorname{DM}(k,\mathbb{Q})$. Moreover, according to [KMP07], this decomposition can be refined: the component $\operatorname{h}^{2}(S)$ can be split into:
@@
representing the algebraic and transcendental parts of $\operatorname{h}^{2}$. Note that by definition, the algebraic part is of the form: $\operatorname{h}^{2}_{alg}=\rho.\mathbbm{1}(-1)[-2]$ where $\rho$ is the Picard number of the surface $S$. Beware that in loc. cit., the motive $\operatorname{h}^{2}_{tr}(S)$ is denoted by $t_{2}(S)$. Recall that it is shown in Corollary 14.8.11 of loc. cit. that the functor $S\mapsto\operatorname{h}^{2}_{tr}(S)$ is a birational invariant of $S$ [^33] As recalled to me by Jan Nagel, this property can also be shown directly by using the blow-up formula for Chow motives combined with the fact any birational morphism of surfaces in characteristic 00 can be written as a sequence of blow-ups of closed points. In particular, it depends only on the function field $K$ of $S$.
Dually, we obtain a decomposition of the homological motive in the homotopy category $\operatorname{DM}(k,\mathbb{Q})$:
@@
If $A$ denotes the Albanese scheme associated with $S$, we have an identification $\operatorname{M}_{1}(S)=\underline{A}$, and by duality: $\operatorname{M}_{3}(S)=\underline{A}^{\vee}(2)[4]\simeq\underline{A}(1)[2]$ (apply 4.3.7). Let us summarize the above discussion.
@@
$\operatorname{M}(S)=\mathbbm{1}\oplus\underline{A}\oplus\rho.\mathbbm{1}(1)[2]\oplus\operatorname{M}_{2}^{tr}(S)\oplus\underline{A}(1)[2]\oplus\mathbbm{1}(2)[4].$
4.4.5.
To go further, we will need to use the coniveau filtration viewed internally within pro-motives, as in [Dég12] or [DP24]. [^34] We follow the notation of the latter which seems better suited. We continue to consider a geometrically connected smooth projective surface $S$ over $k$, with function field $K$.
Recall that a flag in $S$ consists of a decreasing sequence $(Z^{p})_{p\geq 0}$ of closed subschemes of $S$ such that $\operatorname{codim}_{S}(Z^{p})\geq p$. The set $\mathscr{F}lag(S)$ of flags in $S$, ordered by term-wise inclusion, is filtered. We define an increasing filtration of $S$ by pro-schemes, the coniveau filtration, by setting:
$S^{\leq p}:=\underset{Z^{*}\in\mathscr{F}lag(S)}{\text{\rm``lim''}}\big{(}S-Z^{p+1}\big{)}.$
By applying the functor $\operatorname{M}$ component-wise, we then get a canonical homotopy exact sequence in $\operatorname{pro-\mskip-3.0mu}\operatorname{\mathscr{D}M_{gm}}(k)$:
@@
We can then extend the computation of 4.3.2 in the transcendence 2 case as follows.
@@
Consider the above notation. Then there exists homotopy exact sequences in the stable $\infty$-category $\operatorname{pro-\mskip-3.0mu}\operatorname{\mathscr{D}M_{gm}}(k,\mathbb{Q})$:
@@
$\underset{s\in S^{(2)}}{\text{\rm``$\prod$''}}\operatorname{\hat{M}}(\kappa_{s})\{2\}\rightarrow\operatorname{\hat{M}}(S^{\leq 1})\rightarrow\mathbbm{1}\oplus\underline{A}\oplus\rho.\mathbbm{1}(1)[2]\oplus\operatorname{M}_{2}^{tr}(S)\oplus\underline{A}(1)[2]\oplus\mathbbm{1}(2)[4]$
where the $\partial_{x}$ is the residue map (see 3.3.4(D4*)) associated with the valuation on $K$ corresponding to the divisor with generic point $x\in S^{(1)}$.
In particular, the motive of the transcendental degree $2$ function field $K/k$ can be computed in terms of two successive extensions of generic motives associated with fields of lower transcendental degrees on the one hand, and of pure motives, including the transcendental motive of $S$, on the other hand.
@@
This follows immediately from the exact sequence (4.4.5.a) and the preceding discussion, once one notices that $\operatorname{M}(S^{\leq 2})=\operatorname{M}(S)$ and $\operatorname{M}(S^{\leq 0})=\operatorname{M}(S^{=0})=\operatorname{\hat{M}}(K)$. ∎
@@ 5. Motivic cohomology of fields
The structure of the generic motive of a function field $K/k$ is reflected in motivic cohomology, as follows from the formula:
(5.0.1.a) $\operatorname{H}_{\mathrm{M}}^{n,i}(K)=\operatorname{Hom}_{\operatorname{pro-\mskip-3.0mu}\operatorname{\mathscr{D}M}(k)}(\operatorname{\hat{M}}(K),\mathbbm{1}(i)[n]).$
@@
The latter group has several explicit presentations. For the non-trivial case $i>0$, we consider the non-positively graded homological complex made of free abelian groups
@@
where all schemes and products are taken implicitly over the field $K$, $\mathbb{G}_{m}^{\wedge,i}$ is the $i$-th smash product of the pointed scheme $(\mathbb{G}_{m},1)$ (see also 2.2.1 for this notation). Thus the group $c_{K}(\Delta^{n},\mathbb{G}_{m}^{\wedge,i})$ is the free abelian group generated by integral closed subschemes $Z\subset\Delta^{n}\times\mathbb{G}_{m}^{\wedge,i}$ which are finite dominant over $\Delta^{n}_{K}$ and which is not contained in a closed subscheme of the form $\Delta^{n}\times(\mathbb{G}_{m}^{\wedge,j}\times\{1\}\times\mathbb{G}_{m}^{\wedge,i-j-1})$. The differentials are given by the alternating sum of the operation of taking intersection along faces $\Delta^{n-1}_{i}\times\mathbb{G}_{m}^{\wedge,n}$, operations which are always well-defined and can be computed using Serre’s Tor formula.
Thus, the motivic cohomology of the field $K$ can be expressed as follows
@@
As both abelian groups $\operatorname{Ker}(d_{i-n})$ and $\operatorname{Im}(d_{i-n-1})$ are free, this is therefore an explicit presentation of motivic cohomology. While this presentation may not be practical for explicit computations, it provides a bound for the size of motivic cohomology of fields (see 5.2.4 for more). [^35] One gets a similar presentation by considering Bloch’s higher Chow groups.
@@ 5.1. Borel’s computations and $\lambda$-structures
In this subsection we recall the following fundamental computation of motivic cohomology. This is mainly due to Borel if one neglects weights, which were determined by Beilinson using his equally fundamental theory of (higher) regulators. Beilinson’s main argument was explained and corrected in detail by Burgos in [BG02]. We use the notation of 2.2.2 for rational motivic cohomology.
@@
Let $K$ be a number field, and let $r_{1}$ and $r_{2}$ denote the number of real and complex embeddings of $K$, respectively. Then for any pair of integers $(n,i)$, one has:
@@
$\operatorname{H}_{\mathcyr{B}}^{n,i}(\mathcal{O}_{K})=\begin{cases}\mathbb{Q}&n=i=0,\\ \mathcal{O}_{K}^{*}\otimes_{\mathbb{Z}}\mathbb{Q}&n=i=1,\\ \mathbb{Q}^{r_{2}}&n=1,i>1,i\text{ even},\\ \mathbb{Q}^{r_{1}+r_{2}}&n=1,i>1,i\text{ odd},\\ \operatorname{Pic}(\mathcal{O}_{K})_{\mathbb{Q}}&n=2,i=1,\\ 0&\text{otherwise.}\end{cases}$
We will give a proof of this theorem, which is essentially a reduction to Borel’s computation of the stable real cohomology of arithmetic groups ([Bor74]). For clarity, we first recall the arguments in low cohomological degrees, which are now well-established results.
The computations for $i\leq 1$ follow from (2.2.1.b). The same holds for the vanishing $\operatorname{H}_{\mathcyr{B}}^{n,i}(K)$ (resp. $\operatorname{H}_{\mathcyr{B}}^{n,i}(\mathcal{O}_{K})$) for $n-i>0$ (resp. $n-i>1$). We can consider the coniveau spectral sequences for the $1$-dimensional regular scheme $X=\operatorname{Spec}(\mathcal{O}_{K})$ with coefficients in $n$-th twisted rational motivic cohomology and rational K-theory: [^36] The second one is a direct sum of copies of the first one, but it is convenient to use both forms.
@@
$\operatorname{H}_{\mathcyr{B}}^{n,i}(\mathbb{F}_{q})=0$ $\Leftrightarrow(n,i)\neq(0,0),$
$K_{n}(\mathbb{F}_{q})_{\mathbb{Q}}=0$ $\Leftrightarrow n\neq 0.$
@@
Note in particular that $\operatorname{H}_{\mathcyr{B}}^{3,2}(\mathcal{O}_{K})$ must be zero according to the first isomorphism. As a summary, we have obtained the lower $K$-groups as follows:
@@
Thus, it remains to compute $K_{n}(\mathcal{O}_{K})_{\mathbb{Q}}$ for $n\geq 2$, and to determine its graduation for the $\gamma$-filtration to get the above calculation. As already mentioned, the first task was achieved by Borel in his famous paper [Bor74]. The second task follows from the comparison of the Borel and the Beilinson regulators, as detailed in [BG02] (see also [Rap88] for an earlier discussion). In the remainder of this section, we will explicitly recall the computations of Borel and show how to give a direct argument for the determination of the $\gamma$-filtration.
@@
Rational K-theory and homology of groups. [^37] These facts are well-known, but we recall them for completeness, and for introducing our notation. Given any commutative ring $A$ and an integer $n>1$, one gets the following classical computation:
@@
where $\operatorname{GL}(A)$ (resp. $\operatorname{SL}(A)$) is the infinite general (resp. special) linear group associated with $A$, and $X^{+}$ is Quillen $+$-construction applied to a space $X$ [^38] See 6.3.6 for a quick recall in terms of $\infty$-categories. The first isomorphism is Quillen’s comparison of the plus and Q constructions ([Gra76], [God] with more details). To get the second isomorphism, one considers the obvious exact sequence of groups:
@@
After applying the classifying space construction, followed by the plus construction, one gets a fibration sequence in the homotopy category: [^39] Note that the $+$-construction respects fiber sequences, or equivalently homotopy fibers, as it is a localization functor.
$\operatorname{BSL}(A)^{+}\rightarrow\operatorname{BGL}(A)^{+}\rightarrow\mathrm{B}(A^{\times})^{+}.$
Note that $B(A^{\times})^{+}=B(A^{\times})$ since $A^{\times}$ is an abelian group. Thus the associated exact sequence of homotopy groups allows us to conclude.
Next we recall that $\operatorname{BSL}(A)^{+}$ is an H-space. [^40] In fact, the “+=Q” theorem of Quillen shows that it is a loop space. Consequently, the Milnor-Moore/Cartan-Serre theorem implies that the Hurewicz map
$\pi_{n}(\operatorname{BSL}(A)^{+})\rightarrow H_{n}^{\text{sing}}(\operatorname{BSL}(A),\mathbb{Z})$
@@
$H_{*}(SL(A),\mathbb{Q})=H_{*}^{\text{sing}}(\operatorname{BSL}(A),\mathbb{Q})\simeq H_{*}^{\text{sing}}(\operatorname{BSL}(A)^{+},\mathbb{Q}).$
@@
We next fix an integer $r>0$ and consider the arithmetic group $\Gamma_{r}=\operatorname{SL}_{r}(\mathcal{O}_{K})$, which can be seen as a subgroup of the real Lie group [^41] The latter can also be seen as the real points of the algebraic $\mathbb{Q}$-group scheme obtained by the Weil restriction of scalars of the $K$-group scheme $\operatorname{SL}_{r,K}$.
@@
We denote by $\nu_{r}:\Gamma_{r}\rightarrow G_{r}$ the natural inclusion, where $\Gamma_{r}$ is seen with the discrete topology. Viewing $\mathbb{R}$ as a discrete module over $G_{r}$ (resp. $\Gamma_{r}$), one can define the continuous cohomology $H^{*}_{ct}(-,\mathbb{R})$ with coefficient in $\mathbb{R}$ (see e.g. [BW00, IX, §1]). As an application of the main result of [Bor74], one gets the following fundamental computation.